# Personalized Generation In Large Model Era: A Survey

### v1

@@ Abstract
Yiyan Xu1, Jinghao Zhang2, Alireza Salemi3, Xinting Hu4, Wenjie Wang1, Fuli Feng1, Hamed Zamani3, Xiangnan He 1, Tat-Seng Chua5 1University of Science and Technology of China, 2University of Chinese Academy of Sciences 3University of Massachusetts at Amherst, 4Nanyang Technological University 5National University of Singapore {yiyanxu24, wenjiewang96, fulifeng93, xiangnanhe}@gmail.com jinghao.zhang@cripac.ia.ac.cn {asalemi, zamani}@cs.umass.edu xinting001@e.ntu.edu.sg dcscts@nus.edu.sg
@@ 1 Introduction
Despite significant progress Alaluf et al. (2025); Salemi et al. (2024b), research efforts in PGen have largely evolved independently within different communities, such as Natural Language Processing (NLP), Computer Vision (CV), and Information Retrieval (IR). There is no survey that specifically provides a cross-community overview of PGen research. Existing surveys related to PGen separately follow either a model-centric or task-centric perspective, offering only partial summaries of relevant studies. 1) Model-centric surveys focus on specific generative models for personalization, such as Multimodal Large Language Models (MLLMs) Wu et al. (2024b), Large Language Models (LLMs) Zhang et al. (2024j); Chen et al. (2024e); Li et al. (2024i), and Diffusion Models (DMs) Zhang et al. (2024g); 2) Task-centric surveys summarize personalization techniques in specific applications, such as dialogue generation Chen et al. (2024f), role-playing Chen et al. (2024d); Tseng et al. (2024), and generative recommendation Ayemowa et al. (2024). None of these offers a unified framework that comprehensively summarizes PGen research across communities.
@@ 2.3 Workflow
To effectively capture user preferences and specific content needs based on personalized contexts and users’ multimodal instructions, three key techniques are commonly employed: 1) Representation learning, which encodes these inputs into dense embeddings Ruiz et al. (2023); Tang et al. (2024b) or summarizes them into discrete representations (e.g., texts) Shen et al. (2024b); 2) Prompt engineering, which involves designing task-specific prompts to structure user-specific information for generative models Chen et al. (2024g); Li et al. (2025); and 3) Retrieval-augmented generation (RAG), which enriches user-specific information by filtering out irrelevant information and integrating external relevant data Salemi and Zamani (2024); Mysore et al. (2024). By combining these techniques, user modeling establishes a robust foundation for PGen, extracting personalized signals to guide content personalization within the generative modeling process.
@@
  • Step2: Guidance mechanism. To effectively integrate personalized signals, two primary guidance mechanisms are employed: instruction guidance and structural guidance. Specifically, instruction guidance ensures models follow explicit user prompts and instructions using techniques such as in-context learning Xu et al. (2023b); Chen et al. (2024g) and instruction tuning Pi et al. (2024); Xu et al. (2024c). In contrast, structural guidance modifies the model architecture by incorporating additional modules, such as adapters Ye et al. (2023) and cross-attention mechanisms Wei et al. (2023), to embed personalized information.
@@ 3 Personalized Generation Across Modalities
Based on the unified perspective, we present a multi-level taxonomy for PGen, systematically reviewing PGen research across various modalities, personalized contexts, and tasks, as outlined in Table 1. Studies on PGen are first categorized by modality, including text, image, video, audio, 3D, and cross-modal generation. Within each modality, we further classify research based on personalized contexts and examine corresponding tasks and techniques. Additionally, we provide a comprehensive overview of commonly used datasets and evaluation metrics for each modality, which is summarized in Table 2, Table 3, and Table 4.
@@
Personal face/body Face generation PhotoMaker Li et al. (2024k), InstantBooth Shi et al. (2024), InstantID Wang et al. (2024g)
Image (Section 3.2) Virtual try-on IDM-VTON Choi et al. (2025), OOTDiffusion Xu et al. (2024d), OutfitAnyone Sun et al. (2024)
Personalized subjects Subject-driven T2V generation AnimateDiff Guo et al. (2024a), AnimateLCM Wang et al. (2024b), PIA Zhang et al. (2024h)
@@
Personal face Face-to-speech generation VioceMe van Rijn et al. (2022), FR-PSS Wang et al. (2022), Lip2Speech Sheng et al. (2023)
@@
Audio (Section 3.5) Personalized subjects Text-to-audio generation DiffAVA Mo et al. (2023), TAS Li et al. (2024j)
@@
Text (Section 3.1) User Simulation OpinionsQA Santurkar et al. (2023), 3 RoleBench Wang et al. (2024f)
User behaviors General-purpose generation Pinterest Geng et al. (2015), MovieLens Harper and Konstan (2015), MIND Wu et al. (2020b), POG Chen et al. (2019), PASTA Nabati et al. (2024), FABRIC Von Rütte et al. (2023), DialPrompt Liu et al. (2024d), PIP Chen et al. (2024g)
@@
Personalized subjects Subject-driven T2V generation WebVid-10M Bain et al. (2021), UCF101 Soomro (2012), AnimateBench Zhang et al. (2024h), VideoBooth Jiang et al. (2024), StyleCrafter Liu et al. (2024a), Datasets for subject-driven T2I generation…
@@
Audio (Section 3.5) Personalized subjects Text-to-audio generation TASBench Li et al. (2024j), AudioCaps Kim et al. (2019), AudioLDM Liu et al. (2023a)
@@
Face similarity Deng et al. (2019); Schroff et al. (2015); Kim et al. (2022); Wang et al. (2018b) $\checkmark$ Personalization Infinite-ID Wu et al. (2025a), PhotoMaker Li et al. (2024k), ProFusion Zhou et al. (2023)
@@
CLIP-I Radford et al. (2021) $\checkmark$ $\checkmark$ $\checkmark$ Personalization PIA Zhang et al. (2024h), PoseCrafter Zhong et al. (2025), ID-Animator He et al. (2024b)
@@
CLIP-T Radford et al. (2021) $\checkmark$ $\checkmark$ $\checkmark$ Instruction Alignment PIA Zhang et al. (2024h), ConsisID Yuan et al. (2024), PoseCrafter Zhong et al. (2025)
@@
LMD Chen et al. (2018) $\checkmark$ Instruction Alignment DFA-NeRF Yao et al. (2022), DreamTalk Ma et al. (2023), Yi et al. (2020)
@@
Facial Dynamics Devia- tion (FDD) Ma et al. (2024d) $\checkmark$ Personalization DiffSpeaker Ma et al. (2024d), DiffsuionTalker Chen et al. (2023c)
@@
CLIP-T Radford et al. (2021) $\checkmark$ Instruction Alignment MVDream Shi et al. (2023b), DreamBooth3D Raj et al. (2023), MakeYour3D Liu et al. (2025)
@@
Embedding Distance $\checkmark$ $\checkmark$ Personalization UMP Ma et al. (2022), FR-PSS Wang et al. (2022)
@@ 3.1.1 User Behaviors
While recommendation is not directly involved in content generation, it plays a crucial role in delivering personalized content, which has been explored across various scenarios Hou et al. (2024); Harper and Konstan (2015); Wan and McAuley (2018); Wu et al. (2020a). The use of generative models in Recommendation Systems (RecSys) has been widely studied Bao et al. (2023); Wu et al. (2024c); Yang et al. (2023a). Specifically, LLMs have been utilized either through prompting Lyu et al. (2024) or by training them directly to perform recommendation tasks as a form of text generation Lin et al. (2024a). Beyond transformer-based generative models, newer approaches like diffusion models have also been explored for recommendation, highlighting the versatility of generative methods in this domain Wang et al. (2023e); Yang et al. (2024e).
@@ 3.1.2 User Documents
Personalization plays a critical role in enhancing text-based writing assistants, enabling tailored text generation across diverse formats and styles. For this purpose, the LaMP benchmark Salemi et al. (2024b, a); Salemi and Zamani (2024); Zhuang et al. (2024) focuses on short-form text generation, such as creating headlines for news articles or subject lines for emails. In contrast, the LongLaMP benchmark Kumar et al. (2024a) targets longer-form tasks, such as writing product reviews from a user’s perspective based on a rating, generating a post from its summary, or completing an email for a user Salemi et al. (2025b). Additionally, the Personalized Linguistic Attributes Benchmark Alhafni et al. (2024) is designed for text completion tasks, such as extending a post or review, leveraging user-written documents to extract stylistic features that guide LLMs in mimicking a user’s unique writing style. In this domain, RAG is the dominant approach, retrieving relevant information from users’ historically written documents to extract their writing preferences Mysore et al. (2024); Nicolicioiu et al. (2024); Li et al. (2023a).
@@ 3.1.3 User Profiles
In recent years, chatbots and conversational systems have been a central focus of personalized text generation Kottur et al. (2017); Thompson et al. (2004); Kaiss et al. (2023); Kocaballi et al. (2019). The advent of advanced chat models like GPT-4 and Gemini have enabled more sophisticated and personalized interactions. By defining a persona or personality for these models based on user profiles, the system can tailor responses to individual preferences and needs Pradhan and Lazar (2021); Song et al. (2019). To support research in this domain, various dialogue datasets have been developed. For example, LiveChat Gao et al. (2023) is a large-scale dataset created from live streaming interactions, FoCus Jang et al. (2021) focuses on conversational information-seeking scenarios, and Pchatbot Qian et al. (2021) compiles data from Weibo and judicial forums. Enhancing LLMs’ ability to generate personalized responses involves approaches such as zero-shot prompting Zhu et al. (2023b), in-context learning Xu et al. (2023c), and training on limited persona datasets Song et al. (2021); Wang et al. (2024e) or large-scale datasets Zheng et al. (2020); Chen et al. (2023b). Additionally, chain-of-thought reasoning has proven effective in improving alignment with user preferences in dialogue systems Li et al. (2025).
@@ 3.2.1 User Behaviors
This task involves generating tailored images across various scenarios. For instance, PMG Shen et al. (2024b), I-AM-G Wang et al. (2024j), and Pigeon Xu et al. (2024c) utilize historically interacted images of users to infer their visual tastes, enabling personalized generation across various scenarios, including stickers, movie posters, fashion designs, and news posters. In specific, PMG converts historical interacted images into textual descriptions, allowing LLMs to distill user preferences effectively; I-AM-G introduces an interest rewrite strategy to address preference ambiguity and leverages retrieval-augmented generation (RAG) to enrich semantic information; and Pigeon leverages MLLMs with specialized modules to manage noisy historical data and multimodal instructions, ensuring precise user modeling. Similarly, SGDM Xu et al. (2024a) enhances personalization by introducing a style extraction module that captures user-specific style preferences to guide image generation. Personalized PR Chen et al. (2024g) proposes a personalized prompt-rewriting method, leveraging historical user query-image pairs to enhance text-to-image (T2I) personalization.
Beyond inferring user preferences from interaction history, some studies Von Rütte et al. (2023); Liu et al. (2024d); Nabati et al. (2024) have explored interactive, personalized image generation by incorporating real-time user feedback through multi-turn interactions, thereby progressively refining the generated outputs.
@@
This task involves generating personalized fashion designs by inferring personal fashion tastes from user behaviors. Yu et al. (2019) employs GANs to learn user preference vectors from interaction history, generating fashion images compatible with provided images. DiFashion Xu et al. (2024b) adopts DMs to extract user preferences from interaction history for personalized outfit generation and recommendation.
@@
This task aims to create customized, eye-catching visuals for e-commerce products to attract target consumers. Based on user behaviors, AdBooster Shilova et al. (2023) utilizes the Stable Diffusion outpainting model, conditioning on individual user interests to generate appealing product images. Vashishtha et al. (2024) leverages LLMs to generate text prompts for DMs to craft engaging banners. Czapp et al. (2024) employs a contextual bandit algorithm to select prompts from a pool, generating personalized product backgrounds.
@@ 3.2.2 User Profiles
Base on user attributes (e.g., age, gender, interests in characters), LVA-COG Forouzandehmehr et al. (2023) utilizes LLMs to extract user preferences to guide fashion design generation for recommendation.

### v2

@@ Abstract
Yiyan Xu1, Jinghao Zhang2, Alireza Salemi3, Xinting Hu4thanks: Corresponding authors., Wenjie Wang [^1] footnotemark: 1, Fuli Feng1, Hamed Zamani3, Xiangnan He1, Tat-Seng Chua5 1University of Science and Technology of China, 2University of Chinese Academy of Sciences 3University of Massachusetts Amherst, 4Nanyang Technological University 5National University of Singapore {yiyanxu24, wenjiewang96, fulifeng93, xiangnanhe}@gmail.com jinghao.zhang@cripac.ia.ac.cn {asalemi, zamani}@cs.umass.edu xinting001@e.ntu.edu.sg dcscts@nus.edu.sg
@@ 1 Introduction
Despite significant progress Alaluf et al. (2025); Salemi et al. (2024b), research efforts in PGen have largely evolved independently within different communities, such as Natural Language Processing (NLP), Computer Vision (CV), and Information Retrieval (IR). There is no survey that specifically provides a cross-community overview of PGen research. Existing surveys related to PGen separately follow either a model-centric or task-centric perspective, offering only partial summaries of relevant studies. 1) Model-centric surveys focus on specific generative models for personalization, such as Multimodal Large Language Models (MLLMs) Wu et al. (2024b), Large Language Models (LLMs) Zhang et al. (2024l); Chen et al. (2024e); Li et al. (2024j); Liu et al. (2025b); Li et al. (2025a), and Diffusion Models (DMs) Zhang et al. (2024g); 2) Task-centric surveys summarize personalization techniques in specific applications, such as dialogue generation Chen et al. (2024f), role-playing Chen et al. (2024d); Tseng et al. (2024), and generative recommendation Ayemowa et al. (2024). None of these offers a unified framework that comprehensively summarizes PGen research across communities.
@@ 2.3 Workflow
To effectively capture user preferences and specific content needs based on personalized contexts and users’ multimodal instructions, three key techniques are commonly employed: 1) Representation learning, which encodes these inputs into dense embeddings Ruiz et al. (2023); Tang et al. (2024b) or summarizes them into discrete representations (e.g., texts) Shen et al. (2024b); 2) Prompt engineering, which involves designing task-specific prompts to structure user-specific information for generative models Chen et al. (2024g); Li et al. (2025b); and 3) Retrieval-augmented generation (RAG), which enriches user-specific information by filtering out irrelevant information and integrating external relevant data Salemi and Zamani (2024); Mysore et al. (2024). By combining these techniques, user modeling establishes a robust foundation for PGen, extracting personalized signals to guide content personalization within the generative modeling process.
@@
  • Step2: Guidance mechanism. To effectively integrate personalized signals, two primary guidance mechanisms are employed: instruction guidance and structural guidance. Specifically, instruction guidance ensures models follow explicit user prompts and instructions using techniques such as in-context learning Xu et al. (2023b); Chen et al. (2024g); Yang et al. (2023c) and instruction tuning Pi et al. (2024); Xu et al. (2024c). In contrast, structural guidance modifies the model architecture by incorporating additional modules, such as adapters Ye et al. (2023) and cross-attention mechanisms Wei et al. (2023), to embed personalized information.
@@ 3 Personalized Generation Across Modalities
Based on the unified perspective, we present a multi-level taxonomy for PGen, systematically reviewing PGen research across various modalities, personalized contexts, and specific tasks, as outlined in Table 1. Studies on PGen are first categorized by modality, including text, image, video, audio, 3D, and cross-modal generation. Within each modality, we further classify research based on personalized contexts and examine corresponding tasks and techniques. Additionally, we provide a comprehensive overview of commonly used evaluation metrics and datasets for each modality, which is summarized in Table 2, Table 3, and Table 4.
@@
Personal face/body Face generation PhotoMaker Li et al. (2024l), InstantBooth Shi et al. (2024), InstantID Wang et al. (2024g)
Image (Section 3.2) Virtual try-on IDM-VTON Choi et al. (2025), OOTDiffusion Xu et al. (2024d), OutfitAnyone Sun et al. (2024a)
Personalized subjects Subject-driven T2V generation AnimateDiff Guo et al. (2024a), AnimateLCM Wang et al. (2024b), PIA Zhang et al. (2024i)
@@
Personal face Face-to-speech generation VioceMe van Rijn et al. (2022), FR-PSS Wang et al. (2022a), Lip2Speech Sheng et al. (2023)
@@
Audio (Section 3.6) Personalized subjects Text-to-audio generation DiffAVA Mo et al. (2023), TAS Li et al. (2024k)
@@
Text (Section 3.1) User Simulation OpinionsQA Santurkar et al. (2023), 3 RoleBench Wang et al. (2024f), HPD Chen et al. (2023c)
User behaviors General-purpose generation Pinterest Geng et al. (2015), MovieLens Harper and Konstan (2015), MIND Wu et al. (2020b), POG Chen et al. (2019), PASTA Nabati et al. (2024), FABRIC Von Rütte et al. (2023), DialPrompt Liu et al. (2024e), PIP Chen et al. (2024g), U-sticker Chee et al. (2025)
@@
Personalized subjects Subject-driven T2V generation WebVid-10M Bain et al. (2021), UCF101 Soomro (2012), AnimateBench Zhang et al. (2024i), VideoBooth Jiang et al. (2024b), StyleCrafter Liu et al. (2024a), Datasets for subject-driven T2I generation…
@@
Audio (Section 3.6) Personalized subjects Text-to-audio generation TASBench Li et al. (2024k), AudioCaps Kim et al. (2019), AudioLDM Liu et al. (2023a)
@@
Face similarity Deng et al. (2019); Schroff et al. (2015); Kim et al. (2022); Wang et al. (2018b) $\checkmark$ Personalization Infinite-ID Wu et al. (2025a), PhotoMaker Li et al. (2024l), ProFusion Zhou et al. (2023)
@@
CLIP-I Radford et al. (2021) $\checkmark$ $\checkmark$ $\checkmark$ Personalization PIA Zhang et al. (2024i), PoseCrafter Zhong et al. (2025), ID-Animator He et al. (2024b)
@@
CLIP-T Radford et al. (2021) $\checkmark$ $\checkmark$ $\checkmark$ Instruction Alignment PIA Zhang et al. (2024i), ConsisID Yuan et al. (2024), PoseCrafter Zhong et al. (2025)
@@
LMD Chen et al. (2018b) $\checkmark$ Instruction Alignment DFA-NeRF Yao et al. (2022), DreamTalk Ma et al. (2023), Yi et al. (2020)
@@
Facial Dynamics Devia- tion (FDD) Ma et al. (2024d) $\checkmark$ Personalization DiffSpeaker Ma et al. (2024d), DiffsuionTalker Chen et al. (2023d)
@@
CLIP-T Radford et al. (2021) $\checkmark$ Instruction Alignment MVDream Shi et al. (2023b), DreamBooth3D Raj et al. (2023), MakeYour3D Liu et al. (2025a)
@@
Embedding Distance $\checkmark$ $\checkmark$ Personalization UMP Ma et al. (2022), FR-PSS Wang et al. (2022a)
@@ 3.1.1 User Behaviors
While recommendation is not directly involved in content generation, it plays a crucial role in delivering personalized content, which has been explored across various scenarios Hou et al. (2024); Harper and Konstan (2015); Wan and McAuley (2018); Wu et al. (2020a). The use of generative models in Recommendation Systems (RecSys) has been widely studied Bao et al. (2023); Wu et al. (2024c); Yang et al. (2023a); Lin et al. (2025b). Specifically, LLMs have been utilized either through prompting Lyu et al. (2024) or by training them directly to perform recommendation tasks as a form of text generation Lin et al. (2024a). Beyond transformer-based generative models, newer approaches like diffusion models have also been explored for recommendation, highlighting the versatility of generative methods in this domain Wang et al. (2023e); Yang et al. (2024e).
@@ 3.1.2 User Documents
Personalization plays a critical role in enhancing text-based writing assistants, enabling tailored text generation across diverse formats and styles. For this purpose, the LaMP benchmark Salemi et al. (2024b, a); Salemi and Zamani (2024); Zhuang et al. (2024) focuses on short-form text generation, such as creating headlines for news articles or subject lines for emails. In contrast, the LongLaMP benchmark Kumar et al. (2024a) targets longer-form tasks, such as writing product reviews from a user’s perspective based on a rating, generating a post from its summary, or completing an email for a user Salemi et al. (2025b); Liu et al. (2024d); Qiu et al. (2025). Additionally, the Personalized Linguistic Attributes Benchmark Alhafni et al. (2024) is designed for text completion tasks, such as extending a post or review, leveraging user-written documents to extract stylistic features that guide LLMs in mimicking a user’s unique writing style. In this domain, RAG is the dominant approach, retrieving relevant information from users’ historically written documents to extract their writing preferences Mysore et al. (2024); Nicolicioiu et al. (2024); Li et al. (2023a).
@@ 3.1.3 User Profiles
In recent years, chatbots and conversational systems have been a central focus of personalized text generation Kottur et al. (2017); Thompson et al. (2004); Kaiss et al. (2023); Kocaballi et al. (2019). The advent of advanced chat models like GPT-4 and Gemini have enabled more sophisticated and personalized interactions. By defining a persona or personality for these models based on user profiles, the system can tailor responses to individual preferences and needs Pradhan and Lazar (2021); Song et al. (2019). To support research in this domain, various dialogue datasets have been developed. For example, LiveChat Gao et al. (2023) is a large-scale dataset created from live streaming interactions, FoCus Jang et al. (2021) focuses on conversational information-seeking scenarios, and Pchatbot Qian et al. (2021) compiles data from Weibo and judicial forums. Enhancing LLMs’ ability to generate personalized responses involves approaches such as zero-shot prompting Zhu et al. (2023b), in-context learning Xu et al. (2023c), and training on limited persona datasets Song et al. (2021); Wang et al. (2024e) or large-scale datasets Zheng et al. (2020); Chen et al. (2023b). Additionally, chain-of-thought reasoning has proven effective in improving alignment with user preferences in dialogue systems Li et al. (2025b).
@@ 3.2.1 User Behaviors
This task involves generating tailored images across various scenarios. For instance, PMG Shen et al. (2024b), I-AM-G Wang et al. (2024j), Pigeon Xu et al. (2024c), and DRC Xu et al. (2025) utilize historically interacted images of users to infer their visual tastes, enabling personalized generation across various scenarios, including stickers, movie posters, fashion designs, and news posters. In specific, PMG converts historical interacted images into textual descriptions, allowing LLMs to distill user preferences effectively; I-AM-G introduces an interest rewrite strategy to address preference ambiguity and leverages RAG to enrich semantic information; and Pigeon leverages MLLMs with specialized modules to manage noisy historical data and multimodal instructions, ensuring precise user modeling. Similarly, SGDM Xu et al. (2024a) enhances personalization by introducing a style extraction module that captures user-specific style preferences to guide image generation. Personalized PR Chen et al. (2024g) proposes a personalized prompt-rewriting method, leveraging historical user query-image pairs to enhance text-to-image (T2I) personalization.
Beyond inferring user preferences from interaction history, some studies Von Rütte et al. (2023); Liu et al. (2024e); Nabati et al. (2024) have explored interactive, personalized image generation by incorporating real-time user feedback through multi-turn interactions, thereby progressively refining the generated outputs.
@@
This task involves personalized fashion design with inferring personal style preferences from user behaviors. Yu et al. (2019) employs GANs to learn user preference vectors from interaction history, generating fashion images compatible with provided images. DiFashion Xu et al. (2024b) adopts diffusion models to extract user preferences from interaction history for personalized outfit generation and recommendation.
@@
This task aims to create customized, eye-catching visuals for e-commerce products to attract target consumers. Based on user behaviors, AdBooster Shilova et al. (2023) utilizes the Stable Diffusion outpainting model, conditioning on individual user interest to generate appealing product images. Vashishtha et al. (2024) leverages LLMs to generate text prompts for diffusion models to craft engaging banners. Czapp et al. (2024) employs a contextual bandit algorithm to select prompts from a pool, generating personalized product backgrounds.
@@ 3.2.2 User Profiles
Based on user attributes (e.g., age, gender, interests in characters), LVA-COG Forouzandehmehr et al. (2023) utilizes LLMs to extract user preferences to guide fashion design generation. In contrast, PerFusion Lin et al. (2025a) builds a reward model based on user profiles to estimate user preferences, which in turn guides group-level preference optimization of DMs for personalized fashion generation.